OpenStudy (anonymous):

Which polynomial is a perfect square trinomial? 49x^2 - 28x + 16 4a^2 - 20a + 25 25b^2 - 20b - 16 16x^2 - 24x - 9

OpenStudy (anonymous):

second equation

OpenStudy (anonymous):

can you explain why?

OpenStudy (anonymous):

49x2−28x+16 Take the middle term −28 and then divide by two times the first term 2×49, then square it: (−282×49)2=(−27)2=414 This must be equatl to the last term. Since 16≠4/14, this one isn't a perfect square. You have to multiply the first term back http://www.mathsisfun.com/algebra/completing-square.html

OpenStudy (anonymous):

They said d=b2a You want e=0 to have a perfect square which means : c=b24a By the way b24a=a(b2a)2 49x2−28x+16 So b24a=(−28)24(49)=4≠16 4a2−20a+25 We use b24a=(−20)24(4)=40016=25=c So it is a complete square. Do you understand how I put b and a into the equation?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

ANOTHER WAY TO SOLVE b2=4ac Would work, or even b=4ac−−−√ But not ac√=b ...

OpenStudy (anonymous):

THANK YOU!

OpenStudy (anonymous):

yeah :)